OpenStudy (anonymous):

sketch a graph of y=f(x) f(x)=-3x+4

sam (.sam.):

Just sketch it, let y=f(x), then y=-3x+4 ----------------------- when x=1, y=? Find this coordinate Notice that the line intersects the y-axis at (0,4), y=mx+c, and c is the y-intercept Once you have both coordinates, plot it and draw a line

OpenStudy (anonymous):

im lost at graphing @.Sam. im not good at it at all

sam (.sam.):

y=-3x+4 what do you get when x=1, y=?

sam (.sam.):

y=-3(1)+4 y=?

OpenStudy (anonymous):

1@.Sam.

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

 Substitute random values of x in the equation and tabulate y values.